OpenStudy (anonymous):

n a "worst-case" design scenario, a 2000-kg elevator with broken cables is falling at 4.00 m/s when it first contacts a cushioning spring at the bottom of the shaft. The spring is supposed to stop the elevator, compressing 2.00 m as it does so. During the motion a safety clamp applies a constant 17000-N frictional force to the elevator. How would I go about finding the velocity and acceleration of the elevator after it has compressed the spring 1 meter?

OpenStudy (anonymous):

initial kinetic energy of elevator wen touching the cushion is 1/2mv^2 that is converted to potential energy of spring=k*2+frictioanl force*2 find k then using same concept 1/2mv^2=k(1)+frictioanl force*1 v=? 2as=v^2-u^2 v=0 as elevator comes to rest this is all based on moment retriceumptions

OpenStudy (anonymous):

and we are comfortably ignoring gravity........

OpenStudy (anonymous):

to make problem simple

OpenStudy (anonymous):

oh and the spring constant is 1.06 x 10 ^4 N/m. so im confused as to which formula i should use. should i use the 1/2mv^2?

OpenStudy (anonymous):

are u able to follow: 1/2*mv^2=kx+frict force*x v^2-u^2=2ax tahts all see wat energy is converted to what thats all equate forces if equilibrium conserve energy tahts all wat nature and physics is about

OpenStudy (anonymous):

thank you salini

OpenStudy (anonymous):

:)